### v1

# VidHal: Benchmarking Temporal Hallucinations in Vision LLMs

Abstract
Vision Large Language Models (VLLMs) are widely acknowledged to be prone to hallucination. Existing research addressing this problem has primarily been confined to image inputs, with limited exploration of video-based hallucinations. Furthermore, current evaluation methods fail to capture nuanced errors in generated responses, which are often exacerbated by the rich spatiotemporal dynamics of videos. To address this, we introduce VidHal, a benchmark specially designed to evaluate video-based hallucinations in VLLMs. VidHal is constructed by bootstrapping video instances across common temporal aspects. A defining feature of our benchmark lies in the careful creation of captions which represent varying levels of hallucination associated with each video. To enable fine-grained evaluation, we propose a novel caption ordering task requiring VLLMs to rank captions by hallucinatory extent. We conduct extensive experiments on VidHal and comprehensively evaluate a broad selection of models. Our results uncover significant limitations in existing VLLMs regarding hallucination generation. Through our benchmark, we aim to inspire further research on I) holistic understanding of VLLM capabilities, particularly regarding hallucination, and II) extensive development of advanced VLLMs to alleviate this problem. The VidHal dataset and evaluation code are publicly available at https://github.com/Lookuz/VidHal.
1 Introduction
Building on the advancements of Large Language Models (LLMs), Vision LLMs (VLLMs) have recently gained significant attention. Models such as LLaVA [32, 33] have shown impressive performance across various visual understanding tasks involving both images and videos. Despite their potential, VLLMs are notably prone to hallucinations, where generated responses that appear to be plausible contradict the visual context [1]. This problem significantly compromises the reliability of VLLMs, hindering their practical use in real-world applications.
Figure 1: Multiple-Choice Question Answering (MCQA) performance of representative VLLMs on our VidHal benchmark. (Left) Overall ranking of VLLMs. (Right) Detailed accuracy results pertaining to each temporal aspect, wherein higher scores indicate fewer hallucinations.
To tackle this challenge, some methods propose to leverage post-hoc techniques such as contrastive decoding [24, 69, 13, 70] and attention calibration [18, 39, 37, 60, 16, 65, 54]. Other efforts have been devoted to the evaluation of hallucinations in VLLMs. For example, CHAIR [45] initially studies object-based hallucination evaluation with the aid of the image captioning task. Subsequent studies [29, 36, 22, 11] instead harness paired ⟨positive, hallucinatory⟩ questions to probe such hallucinations. Additionally, MMHalBench [48] and AMBER [50] expand beyond object-based evaluations by constructing benchmarks that cover attribute and relationship hallucinations.
Unlike their image-based counterparts, video hallucinations pose unique challenges primarily due to the intricate spatiotemporal dynamics of videos [28, 43, 6, 14, 38, 40]. In particular, video-specific temporal aspects, such as movement direction and chronological order of events, are especially concerning for video-based VLLMs. Furthermore, the richness of video content necessitates a finer-grained understanding, making VLLMs more vulnerable to nuanced hallucinations. Nonetheless, to the best of our knowledge, video-based hallucinations remain underexplored in the existing literature.
To address this research gap, we present VidHal, a benchmark specifically designed to evaluate video-based hallucinations of VLLMs. VidHal features videos that comprehensively cover a wide range of temporal aspects, such as entity actions and sequence of events. Each video is automatically annotated with multiple captions exhibiting varying levels of aspect-specific hallucinations, capturing both subtle and significant discrepancies. In addition, we perform detailed human validation to ensure the robustness and reliability of our annotation process. An additional motivation stems from the limited metrics for quantifying hallucinations in VLLMs. To capture fine-grained hallucinatory errors of these models, we propose a unique caption ordering task that requires models to rank captions by hallucination levels. This consequently leads to a ranking-based NDCG metric and an MCQA accuracy metric, both are distinct from prior ones and specifically tailored to evaluate nuanced hallucinations in video-based VLLMs.
Using our VidHal dataset, we benchmark eleven VLLMs including both open-sourced and proprietary models, with abstracted results summarized in Figure 1. Through these extensive experiments, we identify limitations in nuanced video understanding among all evaluated VLLMs. Specifically, our findings reveal that existing VLLMs struggle to differentiate between captions with varying levels of hallucination. This deficiency is particularly evident when evaluating video-specific aspects, such as Direction and Order, as illustrated in Figure 1, indicating substantial room for improvement in current video-based VLLMs. Besides, we also notice that proprietary models, e.g., GPT-4o [41], frequently outperform open-source models by significant margins.
Overall, the contributions of this work are three-fold:
  • We present VidHal, a benchmark dataset dedicated to video-based hallucination evaluation of VLLMs. Our dataset is distinguished by I) video instances sourced from public video understanding datasets encompassing a diverse range of temporal concepts and II) captions with varying levels of hallucination.
  • We introduce a novel evaluation task of caption ordering along with two metrics designed to evaluate fine-grained hallucination generation in existing VLLMs.
  • We conduct extensive experiments on VidHal with a variety of VLLMs, uncovering limitations in their fine-grained video reasoning abilities, particularly in their tendency to generate hallucinations.
2 Related Work
2.1 Vision Large Language Models
The emergence of powerful LLMs has advanced the development of VLLMs [32, 33, 27, 10, 56, 57, 55]. Typical methods in this category include LLaVA [32], mPLUG-Owl [57, 55, 56], InstructBLIP [10], and MiniGPT-4 [67]. These VLLMs rely on aligning vision encoders with LLMs using connective modules such as Q-Former [10, 26, 27, 61, 9] or MLPs [32, 33, 47] with the instruction tuning stage. Recent methods have extended visual inputs from images to (long) videos, delivering impressive joint spatial-temporal reasoning capabilities. For instance, VideoLLaMA2 [9] enhances the LLaMA model with video understanding capabilities through a Spatial-Temporal Convolution (STC) module. LLaVA-NeXT-Video [34, 62] presents an AnyRes approach that enables reasoning with long videos and high-resolution images.
2.2 Hallucinations in VLLMs
Despite their impressive performance on visual reasoning benchmarks, current VLLMs remain notoriously susceptible to hallucinations [20, 37, 68, 8]. A common demonstration is that the generated responses contain information which is inconsistent with the visual content [1, 35, 59, 53]. Most approaches address the hallucination problem with post-hoc techniques. For example, LURE [66] and Woodpecker [58] develop pipelines that assist VLLMs in revising their responses using expert models. To reduce bias from unimodal and statistical priors, contrastive decoding methods, such as VCD [24] and M3ID [13], along with attention calibration techniques like OPERA [18] are employed to refine token predictions. Building on the success of reinforcement learning for preference optimization [42], HA-DPO [63] and HIO [5] adopt this paradigm to fine-tune VLLMs, yielding outputs with fewer hallucinations. On the other hand, several benchmarks [29, 50, 48, 22, 30, 21, 51, 64, 8, 4, 52, 49, 7] have recently been constructed to quantify visual hallucinations, focusing primarily on object-based hallucinations in images.
Figure 2: Overview of our VidHal benchmark construction pipeline. Using direction as an example from the five selected aspects, we begin by sourcing relevant video instances from existing datasets. Next, the anchor (positive) caption is generated from the original video metadata. Finally, GPT-4o is employed to generate hallucinatory captions at varying levels.
2.3 Video Reasoning Benchmarks
The popularity of video-based VLLMs has been accompanied by the continual development of video benchmarks. Notable ones, such as SEEDBench [25], VideoBench [40], MVBench [28] and VideoMME [14], focus on instances featuring dynamic events that require temporal reasoning beyond a single frame. However, these benchmarks often suffer from a lack of diversity in both reasoning tasks and visual concepts. To address this, AutoEval-Video [6] and Perception Test [43] incorporate a greater variety of complex reasoning tasks, e.g., counterfactual and explanatory reasoning. TempCompass [38] enriches previous benchmarks with a wider range of temporal concepts. Notably, HallusionBench [17], FactVC [31] and VideoCon [2] partially cover video-based hallucinations. However, these benchmarks are limited by their focus on a few video-centric aspects and fail to capture nuanced hallucinations with their evaluation tasks. In contrast, our vidhal is specially designed to carefully evaluate these video-based hallucinations, encompassing a broad range of video aspects with tailored evaluation metrics.
3 VidHal Dataset Construction
We introduce VidHal, a video-language benchmark designed to evaluate hallucinations of Video-LLMs in a holistic manner. As depicted in Figure 2, VidHal comprises of video instances covering various temporal aspects to assess hallucinations across multiple dimensions and granularity of temporal reasoning.
3.1 Temporal Concept Selection
Hallucinations in image inputs can typically be classified into three types: object, attribute, and relation. In contrast, video-based hallucinations in VLLM responses may involve dynamic factors closely associated with the spatiotemporal patterns in video content. In light of this, we propose focusing on the following five aspects to ensure comprehensive coverage of the temporal concepts.
  • Attribute describes the fine-grained characteristics and properties of objects or subjects in the video. We additionally categorize this aspect into sub-aspects of Size, Shape, Colour, Count and State Change.
  • Object relates to the interactions between objects and entities within the video. We further delineate this aspect into two fine-grained sub-aspects: Object Recognition, identifying the objects engaged in interactions, and Interaction Classification, focusing on the semantics of how these objects interact.
  • Action refers to the movements and behaviours exhibited by entities in dynamic video scenes.
  • Direction indicates the orientation and movement trajectory of subjects or objects in the video.
  • Event Order represents the correct sequence of events in the video. During our collection, we retain videos that contain at least three distinct events.
We present one example illustrating the direction aspect in Figure 2, with additional examples available in the supplementary material.
3.2 Hallucinatory Caption Generation
Based on the aspects in Section 3.1, we build our benchmark upon four public video understanding datasets: TempCompass [38], Perception Test [43], MVBench [28] and AutoEval-Video [6]. TempCompass and MVBench provide extensive coverage across all five temporal aspects. Perception Test includes a broader selection of human-object interactions, and AutoEval-Video emphasizes scenarios with attribute changes.
Existing hallucination benchmarks [29, 50] predominantly rely on binary questions for evaluation, which limits their effectiveness for assessing hallucinations in video-based VLLMs. Unlike static images, video content features complex spatiotemporal dynamics that can lead to subtle hallucinations in VLLMs, e.g., minor inconsistencies in events or motion. These nuanced errors, however, are difficult to detect through binary questions.
To address this issue, we advocate a novel evaluation protocol that incorporates several carefully annotated captions. Specifically, each video will be annotated with a set of $M$ captions that reflect varying degrees of hallucination in VLLMs. Given the cost and labor intensity of manual annotation, we opt for automatic caption generation using a carefully designed pipeline illustrated in Figure 2.
Anchor Caption Generation. The video instances in VidHal are sourced from various public datasets, resulting in distinct associated metadata such as long-form captions in AutoEval-Video and question-answer pairs in MVBench and TempCompass. To ensure structure consistency and information granularity in the respective dataset description across all instances, we automatically generate an anchor caption for each video. Specifically, we input the metadata for each video $V^{i}$ into GPT-4o and prompt it to generate a concise and accurate description $y^{i}_{+}$ using the provided metadata information.
Hallucinatory Caption Generation. After obtaining the positive caption for each video instance, we augment the dataset with $M-1$ additional captions containing hallucinated content. For a given video instance $V^{i}$, we construct a set $\mathcal{Y}^{i}_{-}=\{y^{i,1}_{-},\cdots,y^{i,M-1}_{-}\}$ containing captions with different levels of hallucination based on the temporal concepts associated with it. Specifically, $y^{i,k}_{-}$ exhibits heavier hallucination than $y^{i,j}_{-}$ for $j<k$. We leverage GPT-4o to generate $\mathcal{Y}^{i}_{-}$ by combining the anchor caption $y^{i}_{+}$ and prompting it to create $y^{i,1}_{-},\cdots,y^{i,M-1}_{-}$ progressively in increasing levels of hallucination. The set of captions associated with $V^{i}$ is then defined as $\mathcal{Y}^{i}\leftarrow\{y^{i}_{+}\}\bigcup\mathcal{Y}^{i}_{-}$ consisting of both the anchor and hallucinatory captions.
3.3 Dataset Statistics and Human Validation
Our VidHal benchmark consists of a total of 1,000 video instances. Using our automatic annotation pipeline, each video instance is tagged with $M\mathbin{=}3$ captions. As shown in Table 1, our VidHal dataset stands out from other video understanding [25, 40, 28, 14] and hallucination benchmarks [2, 17, 31] in terms of two dimensions: I) VidHalencompasses a diverse range of video-centric temporal aspects; and II) We introduce a novel caption ordering task along with two tailored metrics to capture subtle hallucinations that are previously ignored by paired questions .
Dataset Temporal Aspects Task Formats Evaluation Metrics
Action Attribute Direction Object Order
[…]
Understanding
SEEDBench [25] MCQA Accuracy
VideoBench [40] MCQA Accuracy
MVBench [28] MCQA Accuracy
Video-MME [14] MCQA Accuracy
[…]
Evaluation
FactVC [31] Video Captioning FactVC
VideoCon [2] VLE, NLE ROC-AUC, $Q^{2}$ score
HallusionBench [17] Y/N QA Accuracy
VidHal(Ours) MCQA, Caption Ordering Accuracy, NDCG
Table 1: Comparison of our benchmark dataset with typical video-based datasets. VLE and NLE represent the tasks of visual-language entailment and natural language explanation, respectively.
Figure 3: Human agreement on hallucination levels in the VidHal dataset. (Left) Distribution of agreement ratios per video sample. (Right) Average agreement ratio for each temporal aspect, with an overall average of 87%.
To ensure the reliability of our generated captions at varying levels, we randomly selected 100 examples for human validation, where each sample is labeled by 15 annotators on average. Our human validation process focuses on verifying that the order of hallucinatory captions generated by our pipeline aligns with human judgment. Figure 3 reflects an overall agreement rate of 87% between our automatically generated hallucinatory captions and human annotators, indicating consistency between these two across all video-centric temporal aspects.
4 VidHal Evaluation Protocol
Figure 4: Visual illustration of relative caption ordering task in VidHal. The final ordering is parsed based on VLLM responses for each pair order queried.
Aiming to address the limitations of binary question-based benchmarks, we propose two evaluation tasks: multiple-choice question answering and a novel caption ordering task, detailed in Section 4.1. We also develop corresponding metrics to comprehensively measure hallucinations in video-based VLLMs, elaborated further in Section 4.2.
4.1 Evaluation Tasks
Multiple-Choice Question Answering (MCQA) assesses the model’s spatiotemporal understanding in a coarse-grained manner. Specifically, the model is provided with a video $V^{i}$ and its corresponding set of captions $\mathcal{Y}^{i}$ as answer options. The VLLM is then instructed to select the most appropriate caption for the video.
Caption Ordering evaluates a model’s visual reasoning from a nuanced granularity, instructing VLLMs to order the provided captions based on their hallucination level. Specifically, we design two caption ordering sub-tasks. The first, naive caption ordering, requires VLLMs to rank all captions at once. However, this sub-task can confuse several VLLMs due to its inherently challenging nature and the inferior instruction-following capabilities of some models. As a complement, we propose an additional sub-task, relative caption ordering, which decomposes the prior task into multiple paired caption ordering tasks. Since each paired ordering task is answered in isolation, the VLLM may produce a non-transitive, cyclic ranking. To circumvent this, we first query the model with consecutive caption pairs and prompt it with the final pair only if ambiguities arise in the initial sequence, as illustrated in Figure 4.
Notably, our relative caption ordering task is more challenging than previous binary questions. This complexity arises from certain paired questions in our dataset where both options contain hallucinatory information, making it more difficult to distinguish them apart as opposed to ⟨positive, hallucinatory⟩ pairs.
4.2 Evaluation Metrics
Notations
For a particular video instance $V^{i}$, we define the ground truth caption order for $V^{i}$ to be $\mathcal{Y}^{i}_{*}=(y^{i}_{+},y^{i,1}_{-},\cdots,y^{i,M-1}_{-})$. Further let the $j^{th}$ element in this ordering be indexed as $\mathcal{Y}^{i,j}_{*}$.
MCQA
For the MCQA setting, we adopt the standard accuracy metric,
$\text{Accuracy}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[R_{MCQA}(V^{i},\mathcal{Y}^{i})=y^{i}_{+}\right],$ (1)
where $N$ is the number of video instances, $\mathbb{I}$ denotes the indicator function, and $R_{MCQA}(V^{i},\mathcal{Y}^{i})$ represents the best matched caption from $\mathcal{Y}^{i}$ for $V^{i}$ as predicted by a VLLM.
Caption Ranking
Inspired by metrics from the information retrieval domain [15], we adapt the well-established Normalized Discounted Cumulative Gain (NDCG) [19] for hallucination assessment in VidHal. Unlike previous metrics like POPE [29], our metric awards partial credit for correctly ordered caption pairs even when the optimal ranking is not achieved. As such, we expect the metric to effectively capture both subtle and severe hallucinations generated by video-based VLLMs. Formally, we define our adapted NDCG metric as follows:
$\text{NDCG}=\frac{1}{N}\sum^{N}_{i=1}\frac{\text{DCG}_{i}-\text{rDCG}_{i}}{\text{iDCG}_{i}-\text{rDCG}_{i}},$ (2)
where the DCG is formulated as:
$\text{DCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)},$ (3)
where $\hat{y}^{i,j}$ represents $j^{th}$ caption in the ranked order predicted by the VLLM. The perfect ordering is achieved when $\hat{y}^{i,1}=y^{i}_{+}$ and $\{\hat{y}^{i,j}=y^{i,j-1}_{-}\}_{j=2\to M}$. To evaluate predicted caption orders relative to this ideal sequence, a relevance function $r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)$ is designed to assign higher scores to $\hat{y}^{i,j}$ with lower hallucinatory extent. We formalize the relevance function below:
$r(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})=M+1-\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}),$ (4)
where $\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})$ denotes the position of $\hat{y}^{i,j}$ in $\mathcal{Y}^{i}_{*}$. Finally, $\text{DCG}_{i}$ is normalized to a range of 0 to 1 by introducing $\text{iDCG}_{i}$ and $\text{rDCG}_{i}$. These two terms represent the maximum and minimum $\text{DCG}_{i}$ scores obtained from the optimal ordering $\mathcal{Y}^{i}_{*}$ and its reverse, respectively,
$\text{iDCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\mathcal{Y}^{i,j}_{*},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)},$ (5)
$\text{rDCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\mathcal{Y}^{i,M-j}_{*},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)}.$ (6)
5 Experiments
Model Release Date Visual Encoder LLM #Params #Frames Accuracy NDCG
Naive Relative
Baseline
Random - - - - - 0.326 0.505 0.480
Open-Sourced Models
VideoChat2 [28] Apr. 2024 UMT-L Vicuna 7B 16 0.410 0.490 0.573
Mistral 7B 16 0.524 0.348 0.579
Phi3 3.8B 16 0.468 0.552 0.522
mPLUG-Owl3 [55] Aug. 2024 SigLIP/SO400M Qwen2 7B 16 0.596 0.641 0.707
LLaVA-NeXT-Video [62] Apr. 2024 SigLIP/SO400M Vicuna 7B 32 0.509 0.518 0.620
Qwen1.5 32B 32 0.663 0.641 0.747
VideoLLaMA2 [9] Jun. 2024 CLIP ViT-L/14 Mistral 7B 8 0.541 0.564 0.622
Qwen2 72B 8 0.647 0.787 0.760
Proprietary Models
GPT-4o [41] May 2024 - - 1fps 0.772 0.840 0.826
Gemini-1.5 [44] Feb. 2024 Flash - 1fps 0.657 0.738 0.745
Pro - 1fps 0.671 0.765 0.753
Table 2: Benchmark performance of VLLMs on our VidHal dataset. #Params refers to the number of parameters of the base LLM used. The best performance for each task is highlighted in bold for open-sourced models, and underlined for closed-sourced models.
Figure 5: Aspect-aware results of VLLMs for the (Left) naive and (Right) relative caption ordering task.
5.1 Experimental Settings
Models. We evaluated eleven VLLMs from six different model families, including four open-source models: VideoLLaMA2 [9], VideoChat2 [28], mPLUG-Owl3 [55] and LLaVA-NeXT-Video [62], and two proprietary models: GPT-4o [41] and Gemini-1.5 [44]. These models represent a wide variety of architectural designs and training paradigms. Additionally, we included a random baseline that selects and ranks candidate options randomly.
Implementation Details. All experiments were conducted using four NVIDIA A100 40GB GPUs. The input captions in $\mathcal{Y}^{i}$ are displayed to each model in a randomized order, with a predefined randomization seed applied consistently across all experiments. This operation aims to prevent the model from consistently making the same prediction regardless of the caption orders. We adhered to the inference and model hyperparameters outlined in the respective original models, and employed greedy decoding during generation for a fair comparison.
5.2 Overall Results
Benchmark Results
We present the overall results of representative VLLMs in Table 2 across both MCQA and caption ordering tasks. We make three key observations from this table: I) Proprietary models demonstrate superior results compared to open-sourced models. In particular, GPT-4o achieves the best performance on all tasks, surpassing other models by significant margins. II) Larger VLLMs generally outperform smaller ones in both tasks. This result is supported by the comparison of different LLM bases for the VideoLLaMA2 and LLaVA-NeXT-Video models. III) The caption ordering task poses greater difficulty for current VLLMs than MCQA, evidenced by the larger performance margins between the VLLM models and the random baseline. Notably, the VideoChat2 models show slight to no improvement over the random baseline across both caption ordering tasks. This indicates that current VLLMs greatly suffer from poor fine-grained video understanding and are inclined to generate hallucinations.
Figure 6: Distribution of results of VLLMs across varied input caption orders for the three evaluation tasks.
Aspect-aware Results
Figure 5 highlights the fine-grained, aspect-specific results. It can be observed that these VLLMs exhibit notably stronger performance on action and object aspects than the others. This is likely attributed to current visual instruction tuning data emphasizing object-centric recognition and coarse-grained activity classification, allowing models to leverage image-based priors for predictions. In contrast, these models struggle with fine-grained aspects related to direction and event order, which are especially unique to videos.
5.3 Ablation Studies
Input Order Sensitivity
To assess the robustness of VLLM responses to the order of displayed captions, we conducted additional experiments by evaluating three VLLMs using a fixed static display order across all instances. We repeated this process across all different permutations of input caption order, presenting the results of these models in Figure 6. We observe that the performance of these VLLMs is highly sensitive to the order in which captions are displayed, reflected by their varying results across different order permutations. This instability intensifies with smaller model sizes, with VideoLLaMA2 (7B) showing the highest variance in evaluation results and VideoLLaMA2 (72B) the lowest. Our findings suggest that VLLMs may be particularly vulnerable to input caption order, potentially confounding their performance.
Figure 7: (Top) Averaged consensus score of each respective model, and (Bottom) consensus between each model pairs. Average represents the average consensus score across all VLLMs, with Gemini-1.5 Pro achieving the highest.
Model Correct Incorrect Overall
Naive Caption Ordering
VideoLLaMA2 (7B) 0.889 0.940 0.914
LLaVA-NeXT-Video (32B) 0.727 0.709 0.718
VideoLLaMA2 (72B) 0.726 0.568 0.642
Relative Caption Ordering
VideoLLaMA2 (7B) 0.799 0.781 0.790
LLaVA-NeXT-Video (32B) 0.710 0.579 0.641
VideoLLaMA2 (72B) 0.670 0.444 0.545
Table 3: Overlapping ratios of model predictions under single-frame inputs and full-videos for correctly predicted, incorrectly predicted, and overall of both cases.
Model Consensus
We further assess the reliability of VLLM responses by measuring their consensus on the MCQA task, defined as the proportion of shared correct predictions across correctly answered instances between model pairs. As shown in Figure 7, advanced VLLMs typically show higher consensus with other models sharing similar performance. In contrast, smaller models display less consistency in reasoning, with VideoChat2 demonstrating especially low response agreement among the smaller VLLMs.
Figure 8: (Top) Invalid response rates across all models. VLLMs with no invalid responses are grouped under Others. (Bottom) Regurgitation rates of VLLMs on VidHal. Random and Dataset Statistic indicate the regurgitation rates of the random baseline and ground truth answers, respectively. For both metrics, a lower value indicates better model performance.
Figure 9: Hallucination misalignment (HM) scores on VidHal, with Random representing HM scores from the random baseline.
Figure 10: Qualitative examples of VLLM responses on the caption ordering tasks, for the attribute, order and action aspects.
Image Prior Reliance
Prior research reveals that VLLMs often leverage image priors for reasoning [23, 3], while ignoring the important spatiotemporal features. This phenomenon is exemplified by a few frames having a dominant influence on response generation. To investigate whether the image prior issue influences hallucination generation in video-based VLLMs, we first utilized a video summarization algorithm [46] to extract the most salient frame $v^{i}$ from $V^{i}$. Next, we generated VLLM responses on VidHal using $v^{i}$ instead of $V^{i}$ as the the visual input. The influence of image priors is evaluated by identifying overlapping instances where VLLM responses based on the original video $V^{i}$ and the salient frame $v^{i}$ remain consistent, across both correctly and incorrectly predicted orderings. As illustrated in Table 3, our results highlight that VLLMs exhibit strong reliance on image priors in their predictions. This phenomenon is particularly severe for smaller models such as VideoLLaMA2 (7B).
5.4 In-Depth Analysis
We further leverage three quantitative metrics to identify other deficiencies in VLLM responses.
  • Regurgitation Rate (RR) captures the model’s propensity to consistently generate the same response regardless of input. We define this as the maximum proportion of instances in VidHal where a specific caption order is predicted by the VLLM, across all possible orderings.
  • Hallucination Misalignment (HM) exposes the tendency in visual-language models to favor higher-hallucination captions over lower-hallucination ones in relative caption ranking tasks. We introduce two sub-metrics to detect such misalignments: $HM_{3\rightarrow 1}$ and $HM_{3\rightarrow 2}$, representing the proportion of cases where the VLLM selects the most hallucinatory caption over the anchor and the lower-hallucination caption, respectively. The general form for both terms can be expressed formally as follows:
    $HM_{j\rightarrow k}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[\mathcal{Y}^{i,j}_{*}\prec\mathcal{Y}^{i,k}_{*}\right].$ (7)
  • Invalid Response Rate (IRR) measures the proportion of VLLM responses that fail to provide a valid caption order for the naive caption ordering task.
Figures 8 presents the IRR and RR scores for all evaluated models. Despite being able to formulate responses with the correct structure, many VLLMs tend to produce the same caption order regardless of $V^{i}$, as reflected by their high RR scores. This is particularly evident in models with limited capacity, such as those equipped with LLMs in the 7B parameter range.
Advanced VLLMs, such as VideoLLaMA2 (72B), can generally distinguish positive captions from severely hallucinated ones, as reflected by their low $HM_{3\rightarrow 1}$ scores in Figure 9. However, all evaluated models possess greater $HM_{3\rightarrow 2}$ scores, indicating difficulty in differentiating between two hallucinatory captions with varying levels of hallucination. This gap is especially pronounced in advanced VLLMs like GPT-4o, with $HM_{3\rightarrow 2}$ being significantly closer to the random baseline compared to $HM_{3\rightarrow 1}$. These results reinforce that gaps in nuanced video reasoning may potentially drive hallucinatory behavior in VLLMs, a challenge overlooked by existing paired ⟨positive, hallucinatory⟩-based evaluations [29, 50].
5.5 Qualitative Results
We conducted a qualitative analysis of responses generated by various VLLMs for the caption ordering task, with examples shown in Figure 10. We observe that: I) Relative caption ordering generally guides VLLMs to produce more accurate responses, as evidenced by improvements from naive to relative caption order predictions in most cases. II) Advanced VLLMs exhibit more stable performance across both ordering tasks, reflected by by minimal variation in predictions between the two sub-tasks.
6 Conclusion
In this work, we introduce the VidHal benchmark to address gaps in the video-based hallucination evaluation of VLLMs. VidHal features video instances spanning five temporal aspects. Additionally, we propose a novel caption ordering evaluation task to probe the fine-grained video understanding capabilities of VLLMs. We conduct extensive experiments on VidHal through the evaluation of eleven VLLMs, exposing their limitations in unexpected hallucination generation. Our empirical results shed light on several promising directions for future work: e.g., fine-tuning models with human feedback to reduce undesired regurgitation responses, and countering single-frame priors for improving temporal reasoning. These advancements, on the other hand, will mitigate the hallucination problem in video-based VLLMs, thereby improving their robustness for real-world video understanding applications.
A Benchmark Construction Details
A.1 Dataset Statistics
Figure 11 illustrates the distribution of public dataset sources contributing to the visual instances in VidHal. Additionally, Figures 12 and 13 depict the distribution of temporal aspects across VidHal and the ground truth answers for the MCQA and caption ordering tasks, respectively. One can observe that both temporal aspects and ground truth options are uniformly distributed across our benchmark.
Figure 11: Distribution of visual instances in VidHal by public dataset source, categorized across the five temporal aspects.
Figure 12: Distribution of temporal aspects and subaspects in VidHal.
Figure 13: Distribution of (Left) correct answer options for the MCQA task and (Right) optimal option orders for the caption ordering task.
A.2 Dataset Development Pipeline
Figure 14: Specific skills and corresponding questions from the Perception Test dataset chosen for VidHal instance selection, with the matched aspects indicated in brackets.
Figure 15: Evaluation tasks in MVBench aligned with temporal aspects in VidHal, categorized by aspect.
Visual Instance Selection
To ensure a rich coverage of temporal aspects and visual diversity, we methodically selected video instances from four public datasets: TempCompass [38], Perception Test [43], MVBench [28], and AutoEval Video [6]. Given the unique characteristics of each dataset, we outline the specific guidelines adopted for each dataset below:
  • TempCompass encompasses five temporal aspects: Action, Speed, Direction, Event Order, and Attribute Change. As most of these aspects align with those chosen to construct VidHal, we retain all video instances except those related to speed. TempCompass includes four evaluation tasks: MCQA, Yes/No QA, caption matching, and caption generation. Given the conciseness of captions in the latter two tasks, their information can often be subsumed within the more detailed QA-based annotations. Therefore, we focus exclusively on MCQA and Yes/No QA annotations to create an informative anchor caption.
  • Perception Test spans various skill and reasoning domains to thoroughly evaluate VLLMs’ perception and understanding abilities. Our inspection of these evaluation dimensions reveals alignment between the semantics, physics, and memory skill areas, as well as descriptive and explanatory reasoning dimensions, with the temporal aspects of action, order, and event order. Accordingly, we limit our video selection in Perception Test to these specific pillars. Additionally, we review the question templates adopted in these areas and select video instances with question-answer pairs that support VidHal’s evaluation objectives. The specific skills and associated questions chosen are detailed in Figure 14.
  • MVBench includes twenty video understanding tasks with question-answer pairs designed to challenge the reasoning capabilities of VLLMs. Similar to the Perception Test, we identify the tasks relevant to the temporal aspects in VidHal and focus on collecting videos belonging from these tasks. The specific tasks for each aspect are presented in Figure15. We observe that MVBench contains repeated use of certain scenarios across tasks, indicated by similar question templates. To enhance caption diversity and minimize redundancy, we limit the number of examples for each unique scenario. The collected instances cover all five temporal aspects of VidHal.
  • AutoEval-Video evaluates open-ended response generation in VLLMs through questions with detailed answers across nine skill dimensions. We focus on instances related to the state transition area, specifically assessing changes in object and entity attributes. For each instance, we retain the only answers to associated questions as they act as informative, long-form captions for the video.
Incorrect Anchor Captions
A minority of videos contain anchor captions misaligned with their content, often due to noisy metadata. Such discrepancies subsequently lead to undesirable hallucinatory captions. To remove such instances, we use BLIP2 [27] to calculate frame-text matching scores across all video frames, selecting the maximum score as the representative video-text alignment score. Examples with incorrect anchor captions typically achieve low alignment scores, which are discarded as noisy instances.
LLM-based Caption Generation
We utilize GPT-4o’s [41] text processing and generation capabilities to generate an anchor caption for each selected video, based on metadata from its original public dataset source. This metadata includes QA-based annotations for TempCompass, Perception Test, and MVBench, along with long-form answers for AutoEval-Video. The anchor caption is subsequently used as input for GPT-4o to generate corresponding hallucinatory captions.
To ensure the generated hallucinatory captions meet high-quality standards, we employ a detailed prompt adopting the following strategies to guide GPT-4o’s output:
  • Aspect-specific definitions which outline the characteristics of each aspect to be varied, prompting GPT-4o to modify anchor captions accordingly.
  • Caption construction guidelines that define the structure, format, and hallucination levels required for the generated captions.
  • In-context examples to illustrate the desired form of each hallucinatory caption for each aspect.
The prompts for generating anchor and hallucinatory captions are shown in Figures 16 and 17, respectively. Definitions for each aspect are provided in Figure 18, and aspect-specific in-context examples are detailed in Figures 19 to 27. Separate in-context examples are provided for each Attribute subaspect of Shape, Size, Color, Count, and State Change to account for their distinct natures.
Figure 16: Prompts used for generating the anchor caption from (Top) QA-based annotations and (Bottom) long-form captions.
Figure 17: Prompt for generating aspect-specific hallucinatory captions based on anchor captions and in-context examples.
Figure 18: Definitions incorporated into the prompt for generating hallucinatory captions for each aspect, with separate definitions provided for each sub-aspect in the Attribute aspect.
Figure 19: In-context examples for Size sub-aspect.
Figure 20: In-context examples for Shape sub-aspect.
Figure 21: In-context examples for the Color sub-aspect.
Figure 22: In-context examples for the Count sub-aspect.
Figure 23: In-context examples for the State sub-aspect.
Figure 24: In-context examples for the Object aspect.
Figure 25: In-context examples for the Event Order aspect.
Figure 26: In-context examples for the Action aspect.
Figure 27: In-context examples for the Direction aspect.
Figure 28: Prompt template for evaluating the quality of generated captions for the GPT-4o, Gemini-1.5 Flash, and LLaMA3 (70B) models.
Figure 29: Question prompts for evaluating caption quality based on the three assessment criteria. Prompts with the placeholder <option> are applied individually to the anchor and hallucinatory captions. For question associated with order quality, <option_A> and <option_B> are replaced with the corresponding hallucinatory caption options shown to the LLMs.
Caption Quality Scoring
To identify video instances with the high quality generated captions, we utilize powerful LLMs to evaluate the quality of generated captions. The captions are assessed is based on three specific criteria:
  • Realism determines whether generated scenarios are plausible.
  • Ordering Quality evaluates whether the hallucination level ordering is appropriate.
  • Relevance ensures that deviations from the anchor caption align with the designated aspect.
Binary questions are used to evaluate captions for each criterion, assigning a score of 1 for positive responses, i.e., ”yes”, and 0 otherwise. The scores for each criterion are averaged across all models and prompts, and then summed across all criteria to produce a final quality assessment score for the generated captions of a video instance. We evaluate each set of captions using three LLMs: GPT-4o, Gemini-1.5 Flash [44], and LLaMA3 (70B) [12] along with three variants for each binary question. This ensemble of both models and prompts enhances the robustness of our evaluation.. Figures 28 and 29 provide details of the criterion-specific quality assessment queries and the prompt templates employed for each LLM. We select the top 1,000 examples with the highest quality assessment scores to construct VidHal.
Figure 30: Qualitative examples of video instances and their corresponding generated captions in the VidHal Benchmark, across the five temporal aspects.
A.3 Additional Dataset Examples
We provide additional qualitative examples of video instances and their corresponding captions in Figure 30 for each of the five temporal aspects.
B Human Validation Details
B.1 Human Validation Process
As varying hallucination levels are a distinctive feature of our benchmark, we prioritize validating the robustness of caption ordering produced by our annotation pipeline. Each anchor caption is derived from the original video metadata, making it the most accurate reflection of the video content. Our primary objective is to ensure that the ordering of hallucinatory captions aligns with human judgment. To achieve this, human annotators are shown the video instance along with both hallucinatory captions and are tasked with selecting the caption that better aligns with the video content, as illustrated in Figure 31. Each video instance is reviewed by multiple annotators, with the final human-aligned order determined through a majority vote and compared with our automatically generated order.
Figure 31: Pipeline for validating the quality of generated caption orders in VidHal. For each instance, human annotators are provided with the video and its associated hallucinatory captions. The annotators then select the caption that best aligns with the video content. The selected response is subsequently checked for consistency with the caption with lower hallucination according to our annotation process.
B.2 Misaligned Instances
Table 4 lists video instances that fail to meet the majority agreement threshold established by our annotation process. We additionally provide the corresponding human agreement scores for each instance.
Video ID Agreement Score
action_55 0.429
action_88 0
action_90 0.308
action_118 0.200
action_153 0.250
order_60 0.500
order_109 0.154
attribute_90 0.400
attribute_180 0.071
attribute_192 0.188
object_25 0.375
object_170 0
direction_188 0.400
Table 4: Video examples with generated caption orders misaligned with the human-preferred order during the quality verification process. The agreement score indicates the proportion of human respondents who select the ordering generated in our annotations.
C Evaluation Pipeline Details
C.1 Evaluation Task Prompts
Figures 32 and 33 present the prompts used for the MCQA and naive caption ordering tasks, respectively. The same prompt used for both the MCQA task and the paired questions in the relative caption ordering task.
Figure 32: Prompt template for the MCQA and relative caption ordering evaluation tasks.
Figure 33: Prompt template for the naive caption ordering evaluation task.
C.2 Relative Order Parsing
Prompting the VLLM to predict the order of captions based on their hallucinatory level in the relative caption ordering task involves asking a series of paired questions derived from different caption combinations. However, providing the model with all possible pairs at once may result in cyclic and non-transitive orderings. To address this, we present each caption pair to the VLLM in a systematically selected sequence, beginning with two paired questions. The final paired question is presented to the model to resolve inconsistencies if the multiple possible orderings can be derived from the responses to the first two paired questions. The responses across all paired questions presented to the VLLM is then parsed according to the workflow illustrated in Figure 34.
Figure 34: Decision tree for determining the final caption order based on VLLM responses to paired questions in the relative caption ordering evaluation task.

### v2

@@
 Abstract
-Vision Large Language Models (VLLMs) are widely acknowledged to be prone to hallucination.
+Vision Large Language Models (VLLMs) are widely acknowledged to be prone to hallucinations.
 Existing research addressing this problem has primarily been confined to image inputs, with limited exploration of video-based hallucinations.
@@
 To address this, we introduce VidHal, a benchmark specially designed to evaluate video-based hallucinations in VLLMs.
-VidHal is constructed by bootstrapping video instances across common temporal aspects.
+VidHal is constructed by bootstrapping video instances across a wide range of common temporal aspects.
 A defining feature of our benchmark lies in the careful creation of captions which represent varying levels of hallucination associated with each video.
@@
 Through our benchmark, we aim to inspire further research on I) holistic understanding of VLLM capabilities, particularly regarding hallucination, and II) extensive development of advanced VLLMs to alleviate this problem.
-The VidHal dataset and evaluation code are publicly available at https://github.com/Lookuz/VidHal.
+Our VidHal dataset and evaluation code are publicly available at https://github.com/Lookuz/VidHal.
 1 Introduction
@@
 Models such as LLaVA [33, 34] have shown impressive performance across various visual understanding tasks involving both images and videos.
-Despite their potential, VLLMs are notably prone to hallucinations, where generated responses that appear to be plausible contradict the visual context [1].
+Despite their potential, VLLMs are notably prone to hallucinations, where generated responses that appear to be plausible contradict the visual context [1, 57].
 This problem significantly compromises the reliability of VLLMs, hindering their practical use in real-world applications.
@@
 To address this research gap, we present VidHal, a benchmark specifically designed to evaluate video-based hallucinations of VLLMs.
-VidHal features videos that comprehensively cover a wide range of temporal aspects, such as entity actions and sequence of events.
+VidHal features videos that comprehensively cover a broad range of temporal aspects, such as entity actions and sequence of events.
 Each video is automatically annotated with multiple captions exhibiting varying levels of aspect-specific hallucinations, capturing both subtle and significant discrepancies.
@@
 This consequently leads to a ranking-based NDCG metric and an MCQA accuracy metric, both are distinct from prior ones and specifically tailored to evaluate nuanced hallucinations in video-based VLLMs.
-Using our VidHal dataset, we benchmark eleven VLLMs including both open-sourced and proprietary models, with abstracted results summarized in Figure 1.
+Using our VidHal dataset, we benchmark thirteen VLLMs including both open-sourced and proprietary models, with abstracted results summarized in Figure 1.
 Through these extensive experiments, we identify limitations in nuanced video understanding among all evaluated VLLMs.
@@
 For instance, VideoLLaMA2 [8] enhances the LLaMA model with video understanding capabilities through a Spatial-Temporal Convolution (STC) module.
-LLaVA-NeXT-Video [34, 62] presents an AnyRes approach that enables reasoning with long videos and high-resolution images.
+LLaVA-NeXT-Video [35, 66] presents an AnyRes approach that enables reasoning with long videos.
 2.2 Hallucinations in VLLMs
@@
 To reduce bias from unimodal and statistical priors, contrastive decoding methods, such as VCD [23] and M3ID [12], along with attention calibration techniques like OPERA [17] are employed to refine token predictions.
-Building on the success of reinforcement learning for preference optimization [42], HA-DPO [63] and HIO [5] adopt this paradigm to fine-tune VLLMs, yielding outputs with fewer hallucinations.
-On the other hand, several benchmarks [29, 50, 48, 22, 30, 21, 51, 64, 8, 4, 52, 49, 7] have recently been constructed to quantify visual hallucinations, focusing primarily on object-based hallucinations in images.
+Building on the success of reinforcement learning for preference optimization in LLM development [43], HA-DPO [67], POVID [70] and CSR [72] adopt this paradigm to fine-tune VLLMs, yielding outputs with fewer hallucinations.
 Figure 2: Overview of our VidHal benchmark construction pipeline.
@@
 2.3 Video Reasoning Benchmarks
-The popularity of video-based VLLMs has been accompanied by the continual development of video benchmarks.
-Notable ones, such as SEEDBench [25], VideoBench [40], MVBench [28] and VideoMME [14], focus on instances featuring dynamic events that require temporal reasoning beyond a single frame.
-However, these benchmarks often suffer from a lack of diversity in both reasoning tasks and visual concepts.
-To address this, AutoEval-Video [6] and Perception Test [43] incorporate a greater variety of complex reasoning tasks, e.g., counterfactual and explanatory reasoning.
-TempCompass [38] enriches previous benchmarks with a wider range of temporal concepts.
-Notably, HallusionBench [17], FactVC [31] and VideoCon [2] partially cover video-based hallucinations.
-However, these benchmarks are limited by their focus on a few video-centric aspects and fail to capture nuanced hallucinations with their evaluation tasks.
-In contrast, our vidhal is specially designed to carefully evaluate these video-based hallucinations, encompassing a broad range of video aspects with tailored evaluation metrics.
+The rise of video-based VLLMs has driven the development of numerous video benchmarks.
+Notable examples, such as SEEDBench [24], VideoBench [41], MVBench [28], and VideoMME [13], focus on dynamic events requiring temporal reasoning beyond individual frames.
+However, these benchmarks often lack diversity in reasoning tasks and visual concepts.
+To address this, AutoEval-Video [5] and Perception Test [44] introduce complex reasoning tasks such as counterfactual and explanatory reasoning, while TempCompass [39] expands temporal concept coverage.
+Several benchmarks [29, 51, 49, 21, 31, 20, 53, 68, 7, 4, 54, 50, 6] have been constructed to quantify visual hallucinations, primarily targeting object-based hallucinations in images.
+HallusionBench [16], VideoCon [2], and Vript [58] provides partial coverage of video-based hallucinations, while VideoHallucer [52] introduces a benchmark for hallucination detection in videos.
+However, these benchmarks provide limited coverage of spatio-temporal concepts, focusing on conventional aspects like actions and objects while neglecting other video-centric elements such as direction.
+Additionally, their evaluation strategies primarily follow image-based approaches, which we argue are less effective in capturing nuanced, video-specific hallucinations.
 3 VidHal Dataset Construction
-We introduce VidHal, a video-language benchmark designed to evaluate hallucinations of Video-LLMs in a holistic manner.
-As depicted in Figure 2, VidHal comprises of video instances covering various temporal aspects to assess hallucinations across multiple dimensions and granularity of temporal reasoning.
+We introduce VidHal, a unique video-language benchmark designed to evaluate hallucinations of Video-LLMs in a comprehensive manner.
+As depicted in Figure 2, VidHal comprises of video instances which span a diverse spectrum of temporal aspects, including previously unexplored aspects such as directional movement.
+In contrast to previous studies on video hallucination evaluation [58, 52, 2], VidHal incorporates multiple hallucinated captions per video, enabling the assessment of video hallucinations at multiple levels of granularity.
+3.1 Temporal Hallucinations in Videos
+Hallucinations in VLLMs occur when the model fabricates details in its responses that contradict the provided visual content.
+Compared to images, video hallucinations extend beyond static visual elements to include misperceptions of dynamic changes within scenes.
+We categorize these temporal hallucinations into two semantic levels:
+Lexical Semantics (L-Sem) captures instances where VLLMs misinterpret words related to temporal features, including nouns referring to objects or attributes (e.g., misidentifying a color change from green to red as green to orange) and verbs describing actions (e.g., interpreting ”kicking a ball” as ”throwing a ball”).
+Clause Semantics (C-Sem) encompasses errors involving event descriptions and their sequences, where the VLLM incorrectly predicts the order of events occurring in the video.
+For example, given sequentially occurring events $A$ and $B$ in a video, the model may perceive $B$ preceding $A$.
+By addressing these two dimensions of video-based hallucinations, VidHal offers holistic coverage over the level of detail in which VLLMs may hallucinate.
 3.2 Temporal Concept Selection
-Hallucinations in image inputs can typically be classified into three types: object, attribute, and relation.
-In contrast, video-based hallucinations in VLLM responses may involve dynamic factors closely associated with the spatiotemporal patterns in video content.
-In light of this, we propose focusing on the following five aspects to ensure comprehensive coverage of the temporal concepts.
-  • Attribute describes the fine-grained characteristics and properties of objects or subjects in the video.
-We additionally categorize this aspect into sub-aspects of Size, Shape, Colour, Count and State Change.
-  • Object relates to the interactions between objects and entities within the video.
-We further delineate this aspect into two fine-grained sub-aspects: Object Recognition, identifying the objects engaged in interactions, and Interaction Classification, focusing on the semantics of how these objects interact.
-  • Action refers to the movements and behaviours exhibited by entities in dynamic video scenes.
-  • Direction indicates the orientation and movement trajectory of subjects or objects in the video.
-  • Event Order represents the correct sequence of events in the video. During our collection, we retain videos that contain at least three distinct events.
-We present one example illustrating the direction aspect in Figure 2, with additional examples available in the supplementary material.
+Prior research on hallucination evaluation for both images [29, 51, 46] and videos [52, 58, 16] has predominantly focused on common visual aspects such as action- and object-based hallucinations.
+However, video-based hallucinations may involve additional dynamic factors associated with spatio-temporal patterns, which are overlooked in these studies.
+In light of this, we propose to focus on the following five aspects to ensure comprehensive coverage of temporal concepts.
+Specifically, the first four aspects address hallucinations based on lexical semantics, while the fifth targets those based on clause-level semantics.
+  • Attribute (L-Sem) describes the fine-grained characteristics and properties of objects or subjects in the video.
+We additionally categorize this aspect into sub-aspects of Size, Shape, Color, Count and State Change.
+  • Object (L-Sem) relates to the interactions between objects and entities within the video.
+We further delineate this aspect into two fine-grained sub-aspects: Object Recognition, identifying the objects engaged in interactions, and Interaction Classification which concentrate on how these objects interact with other objects or subjects.
+  • Action (L-Sem) refers to the movements and behaviours exhibited by entities in dynamic video scenes.
+  • Direction (L-Sem) indicates the orientation and movement trajectory of subjects or objects in the video.
+  • Event Order (C-Sem) represents the correct sequence of events in the video. During our collection, we retain videos that contain at least three distinct events.
+We present an example that illustrates the direction aspect in Figure 2, with additional examples available in the supplementary material.
 3.3 Hallucinatory Caption Generation
 Based on the aspects in Section 3.2, we build our benchmark upon four public video understanding datasets: TempCompass [39], Perception Test [44], MVBench [28] and AutoEval-Video [5].
-TempCompass and MVBench provide extensive coverage across all five temporal aspects.
-Perception Test includes a broader selection of human-object interactions, and AutoEval-Video emphasizes scenarios with attribute changes.
-Existing hallucination benchmarks [29, 50] predominantly rely on binary questions for evaluation, which limits their effectiveness for assessing hallucinations in video-based VLLMs.
-Unlike static images, video content features complex spatiotemporal dynamics that can lead to subtle hallucinations in VLLMs, e.g., minor inconsistencies in events or motion.
-These nuanced errors, however, are difficult to detect through binary questions.
-To address this issue, we advocate a novel evaluation protocol that incorporates several carefully annotated captions.
+TempCompass and MVBench extensively cover all five temporal aspects, while Perception Test and AutoEval-Video highlights human-object interactions and attribute changes, respectively.
+Existing hallucination benchmarks [29, 51] rely mostly on binary questions for evaluation, limiting their efficacy in detecting subtle video hallucinations, such as minor event inconsistencies.
+To address this issue, we advocate a novel evaluation protocol incorporating several carefully annotated captions.
 Specifically, each video will be annotated with a set of $M$ captions that reflect varying degrees of hallucination in VLLMs.
@@
 Anchor Caption Generation.
-The video instances in VidHal are sourced from various public datasets, resulting in distinct associated metadata such as long-form captions in AutoEval-Video and question-answer pairs in MVBench and TempCompass.
+The video instances in VidHal are sourced from various public datasets, resulting in distinct associated metadata such as long-form captions in AutoEval-Video and question-answer pairs in MVBench.
 To ensure structure consistency and information granularity in the respective dataset description across all instances, we automatically generate an anchor caption for each video.
@@
 The set of captions associated with $V^{i}$ is then defined as $\mathcal{Y}^{i}\leftarrow\{y^{i}_{+}\}\bigcup\mathcal{Y}^{i}_{-}$ consisting of both the anchor and hallucinatory captions.
+Figure 3: Human agreement on hallucination levels in the VidHal dataset.
+(Left) Distribution of agreement ratios per video sample.
+(Right) Average agreement ratio for each temporal aspect, with an overall average of 87%.
 3.4 Dataset Statistics and Human Validation
@@
 Action Attribute Direction Object Order
+Size Shape Color Count State-Change Recognition Interaction
 […]
-Understanding
+Reasoning
 SEEDBench [24] MCQA Accuracy
@@
 Evaluation
-FactVC [31] Video Captioning FactVC
-VideoCon [2] VLE, NLE ROC-AUC, $Q^{2}$ score
+Vript [58] Video Captioning F1 Score
+Event Ordering Accuracy
+VideoCon [2] VL Entailment ROC-AUC
 HallusionBench [16] Y/N QA Accuracy
-VidHal(Ours) MCQA, Caption Ordering Accuracy, NDCG
-Table 1: Comparison of our benchmark dataset with typical video-based datasets. VLE and NLE represent the tasks of visual-language entailment and natural language explanation, respectively.
-Figure 3: Human agreement on hallucination levels in the VidHal dataset.
-(Left) Distribution of agreement ratios per video sample.
-(Right) Average agreement ratio for each temporal aspect, with an overall average of 87%.
+VidHal(Ours) MCQA Accuracy
+Caption Ordering NDCG
+Table 1: Comparison of our benchmark dataset with existing video-based reasoning and hallucination evaluation datasets.
+For datasets with multiple evaluation tasks, only those relevant to hallucination evaluation are included.
+VL Entailment denotes the task of video-language entailment, while Event Ordering prompts the model to determine the chronological sequence of scenes in a video.
 To ensure the reliability of our generated captions at varying levels, we randomly selected 100 examples for human validation, where each sample is labeled by 15 annotators on average.
@@
 Notably, our relative caption ordering task is more challenging than previous binary questions.
-This complexity arises from certain paired questions in our dataset where both options contain hallucinatory information, making it more difficult to distinguish them apart as opposed to ⟨positive, hallucinatory⟩ pairs.
+This complexity arises from certain paired questions in VidHal where both options are hallucinatory, making them harder to distinguish as opposed to ⟨positive, hallucinatory⟩ pairs.
+Model Release Date Vision Encoder LLM #Params #Frames Accuracy NDCG
+Naive Relative
+Baseline
+Random - - - - - 0.326 0.505 0.480
+Open-Sourced Models
+VideoChat [27] May 2023 EVA-CLIP-G Vicuna 7B 8 0.381 0.475 0.488
+LLaMA-VID [30] Nov. 2023 EVA-CLIP-G Vicuna 7B 1fps 0.358 0.486 0.521
+VideoChat2 (Vicuna) [28] Nov. 2023 UMT-L Vicuna 7B 16 0.426 0.486 0.577
+VideoChat2 (Mistral) May 2024 UMT-L Mistral 7B 16 0.443 0.503 0.475
+VideoChat2 (Phi) Jun. 2024 UMT-L Phi3 3.8B 16 0.514 0.626 0.612
+mPLUG-Owl3 [59] Aug. 2024 SigLIP/SO400M Qwen2 7B 16 0.596 0.641 0.707
+LLaVA-NeXT-Video (7B) [66] Apr. 2024 SigLIP/SO400M Vicuna 7B 32 0.509 0.518 0.620
+LLaVA-NeXT-Video (32B) Apr. 2024 SigLIP/SO400M Qwen1.5 32B 32 0.663 0.641 0.747
+VideoLLaMA2 (7B) [8] Jun. 2024 CLIP ViT-L/14 Mistral 7B 8 0.541 0.564 0.622
+VideoLLaMA2 (72B) Jun. 2024 CLIP ViT-L/14 Qwen2 72B 8 0.647 0.787 0.760
+Proprietary Models
+GPT-4o [42] May 2024 - - - 1fps 0.772 0.840 0.826
+Gemini-1.5 (Flash) [45] May 2024 - - - 1fps 0.657 0.738 0.745
+Gemini-1.5 (Pro) May 2024 - - - 1fps 0.671 0.765 0.753
+Table 2: Benchmark performance of VLLMs on our VidHal dataset.
+#Params refers to the number of parameters of the base LLM used.
+The best performance for each task is highlighted in bold for open-sourced models, and underlined for closed-sourced models.
 4.2 Evaluation Metrics
@@
 MCQA
-For the MCQA setting, we adopt the standard accuracy metric,
+We employ the standard accuracy metric:
 $\text{Accuracy}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[R_{MCQA}(V^{i},\mathcal{Y}^{i})=y^{i}_{+}\right],$ (1)
@@
 Unlike previous metrics like POPE [29], our metric awards partial credit for correctly ordered caption pairs even when the optimal ranking is not achieved.
-As such, we expect the metric to effectively capture both subtle and severe hallucinations generated by video-based VLLMs.
+As such, we expect the metric to effectively capture and distinguish both subtle and severe hallucinations generated by video-based VLLMs.
 Formally, we define our adapted NDCG metric as follows:
 $\text{NDCG}=\frac{1}{N}\sum^{N}_{i=1}\frac{\text{DCG}_{i}-\text{rDCG}_{i}}{\text{iDCG}_{i}-\text{rDCG}_{i}},$ (2)
-where the DCG is formulated as:
+where $\text{DCG}_{i}$ is formulated as:
 $\text{DCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)},$ (3)
-where $\hat{y}^{i,j}$ represents $j^{th}$ caption in the ranked order predicted by the VLLM.
+and $\hat{y}^{i,j}$ represents $j^{th}$ caption in the ranked order predicted by the VLLM.
 The perfect ordering is achieved when $\hat{y}^{i,1}=y^{i}_{+}$ and $\{\hat{y}^{i,j}=y^{i,j-1}_{-}\}_{j=2\to M}$.
 To evaluate predicted caption orders relative to this ideal sequence, a relevance function $r\left(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}\right)$ is designed to assign higher scores to $\hat{y}^{i,j}$ with lower hallucinatory extent.
-We formalize the relevance function below:
 $r(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})=M+1-\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*}),$ (4)
 where $\text{pos}(\hat{y}^{i,j},\mathcal{Y}^{i}_{*})$ denotes the position of $\hat{y}^{i,j}$ in $\mathcal{Y}^{i}_{*}$.
-Finally, $\text{DCG}_{i}$ is normalized to a range of 0 to 1 by introducing $\text{iDCG}_{i}$ and $\text{rDCG}_{i}$.
-These two terms represent the maximum and minimum $\text{DCG}_{i}$ scores obtained from the optimal ordering $\mathcal{Y}^{i}_{*}$ and its reverse, respectively,
+Finally, $\text{DCG}_{i}$ is normalized to a range of $[0,1]$ using $\text{iDCG}_{i}$ and $\text{rDCG}_{i}$, with a score of 1 indicating perfect alignment of the predicted order with $\mathcal{Y}^{i}_{*}$.
+Specifically, these terms represent the maximum and minimum $\text{DCG}_{i}$ scores obtained from the optimal ordering $\mathcal{Y}^{i}_{*}$ and its reverse, respectively,
 $\text{iDCG}_{i}=\sum^{M}_{j=1}\frac{r\left(\mathcal{Y}^{i,j}_{*},\mathcal{Y}^{i}_{*}\right)}{\log(j+1)},$ (5)
@@
 5 Experiments
-Model Release Date Visual Encoder LLM #Params #Frames Accuracy NDCG
-Naive Relative
-Baseline
-Random - - - - - 0.326 0.505 0.480
-Open-Sourced Models
-VideoChat2 [28] Apr. 2024 UMT-L Vicuna 7B 16 0.410 0.490 0.573
-Mistral 7B 16 0.524 0.348 0.579
-Phi3 3.8B 16 0.468 0.552 0.522
-mPLUG-Owl3 [55] Aug. 2024 SigLIP/SO400M Qwen2 7B 16 0.596 0.641 0.707
-LLaVA-NeXT-Video [62] Apr. 2024 SigLIP/SO400M Vicuna 7B 32 0.509 0.518 0.620
-Qwen1.5 32B 32 0.663 0.641 0.747
-VideoLLaMA2 [9] Jun. 2024 CLIP ViT-L/14 Mistral 7B 8 0.541 0.564 0.622
-Qwen2 72B 8 0.647 0.787 0.760
-Proprietary Models
-GPT-4o [41] May 2024 - - 1fps 0.772 0.840 0.826
-Gemini-1.5 [44] Feb. 2024 Flash - 1fps 0.657 0.738 0.745
-Pro - 1fps 0.671 0.765 0.753
-Table 2: Benchmark performance of VLLMs on our VidHal dataset.
-#Params refers to the number of parameters of the base LLM used.
-The best performance for each task is highlighted in bold for open-sourced models, and underlined for closed-sourced models.
-Figure 5: Aspect-aware results of VLLMs for the (Left) naive and (Right) relative caption ordering task.
 5.1 Experimental Settings
 Models.
-We evaluated eleven VLLMs from six different model families, including four open-source models: VideoLLaMA2 [9], VideoChat2 [28], mPLUG-Owl3 [55] and LLaVA-NeXT-Video [62], and two proprietary models: GPT-4o [41] and Gemini-1.5 [44].
+We evaluated thirteen VLLMs from eight different model families, including six open-source models: VideoChat [27], LLaMA-VID [30], VideoLLaMA2 [8], VideoChat2 [28], mPLUG-Owl3 [59] and LLaVA-NeXT-Video [66], and two proprietary models: GPT-4o [42] and Gemini-1.5 [45].
 These models represent a wide variety of architectural designs and training paradigms.
@@
 All experiments were conducted using four NVIDIA A100 40GB GPUs.
-The input captions in $\mathcal{Y}^{i}$ are displayed to each model in a randomized order, with a predefined randomization seed applied consistently across all experiments.
-This operation aims to prevent the model from consistently making the same prediction regardless of the caption orders.
+The input captions in $\mathcal{Y}^{i}$ were presented in a randomized order using a fixed, predefined randomization seed across experiments.
 We adhered to the inference and model hyperparameters outlined in the respective original models, and employed greedy decoding during generation for a fair comparison.
@@
 III) The caption ordering task poses greater difficulty for current VLLMs than MCQA, evidenced by the larger performance margins between the VLLM models and the random baseline.
-Notably, the VideoChat2 models show slight to no improvement over the random baseline across both caption ordering tasks.
+Notably, VideoChat and VideoChat2 (Mistral) show slight to no improvement over the random baseline across both caption ordering tasks.
 This indicates that current VLLMs greatly suffer from poor fine-grained video understanding and are inclined to generate hallucinations.
-Figure 6: Distribution of results of VLLMs across varied input caption orders for the three evaluation tasks.
+Figure 5: Aspect-aware results of VLLMs for the (Left) naive and (Right) relative caption ordering task. The dotted lines represent the average NDCG scores across all models.
 Aspect-aware Results
@@
 In contrast, these models struggle with fine-grained aspects related to direction and event order, which are especially unique to videos.
+We further analyzed the distribution of results for the relative caption ranking task across sub-aspects of the Attribute and Object aspects in Figure 6.
+While VLLMs generally maintain consistent performance across Attribute sub-aspects, they demonstrated slightly less effectiveness in inference involving Count and Color.
+For the Object aspect, several models performed significantly worse in Interaction Classification than in Object Recognition.
+This highlights the need for greater emphasis on modeling object-based interactions to bridge the gap between object recognition and interaction understanding in VLLMs.
+Figure 6: NDCG scores for the Attribute (Left) and Object (Right) sub-aspects in the relative caption ranking task.
 5.3 Ablation Studies
-Input Order Sensitivity
-To assess the robustness of VLLM responses to the order of displayed captions, we conducted additional experiments by evaluating three VLLMs using a fixed static display order across all instances.
-We repeated this process across all different permutations of input caption order, presenting the results of these models in Figure 6.
-We observe that the performance of these VLLMs is highly sensitive to the order in which captions are displayed, reflected by their varying results across different order permutations.
-This instability intensifies with smaller model sizes, with VideoLLaMA2 (7B) showing the highest variance in evaluation results and VideoLLaMA2 (72B) the lowest.
-Our findings suggest that VLLMs may be particularly vulnerable to input caption order, potentially confounding their performance.
-Figure 7: (Top) Averaged consensus score of each respective model, and (Bottom) consensus between each model pairs.
-Average represents the average consensus score across all VLLMs, with Gemini-1.5 Pro achieving the highest.
-Model Correct Incorrect Overall
-Naive Caption Ordering
-VideoLLaMA2 (7B) 0.889 0.940 0.914
-LLaVA-NeXT-Video (32B) 0.727 0.709 0.718
-VideoLLaMA2 (72B) 0.726 0.568 0.642
-Relative Caption Ordering
-VideoLLaMA2 (7B) 0.799 0.781 0.790
-LLaVA-NeXT-Video (32B) 0.710 0.579 0.641
-VideoLLaMA2 (72B) 0.670 0.444 0.545
-Table 3: Overlapping ratios of model predictions under single-frame inputs and full-videos for correctly predicted, incorrectly predicted, and overall of both cases.
-Model Consensus
-We further assess the reliability of VLLM responses by measuring their consensus on the MCQA task, defined as the proportion of shared correct predictions across correctly answered instances between model pairs.
-As shown in Figure 7, advanced VLLMs typically show higher consensus with other models sharing similar performance.
-In contrast, smaller models display less consistency in reasoning, with VideoChat2 demonstrating especially low response agreement among the smaller VLLMs.
+VideoChat2 (Vicuna) mPLUG-Owl3 LLaVA-NeXT -Video (7B) VideoLLaMA2 (7B)
+$P_{\text{TI}}\ (\%)$ 7.3 3.0 3.7 4.0
+Table 3: Percentage of VidHal instances where final predicted orders contradict individual paired-caption orders.
+Transitive Robustness
+Despite improving VLLM stability in caption ordering, the relative caption ranking task may reveal transitive inconsistencies between the final caption order and individual paired orders.
+For instance, the predicted order $\mathcal{\hat{Y}}^{i}=[\hat{y}^{i,1},\hat{y}^{i,2},\hat{y}^{i,3}]$ may be inferred when the VLLM predicts $\hat{y}^{i,1}\prec\hat{y}^{i,2}$ and $\hat{y}^{i,2}\prec\hat{y}^{i,3}$, but it may predict $\hat{y}^{i,3}\prec\hat{y}^{i,1}$ when prompted to order these two captions, contradicting the initial order.
+We quantify such errors by measuring the proportion of instances where $\mathcal{\hat{Y}}^{i}$ does not entail $\hat{y}^{i,3}\prec\hat{y}^{i,1}$.
+This is formally defined as:
+$P_{\text{TI}}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[R_{MCQA}(V^{i},\{\hat{y}^{i,1},\hat{y}^{i,3}\})=\hat{y}^{i,3}\right].$ (7)
+The results, presented in Table 3, demonstrate strong consistency between individual paired caption orders and final caption orders in VLLM predictions.
+Figure 7: Overlapping ratios of model predictions under single-frame and full-video inputs for correct, incorrect and overall predictions.
+Complete Reliance indicates that the VLLM always produces the same response for both video and single-frame inputs.
+Image Prior Reliance
+Prior research reveals that VLLMs often leverage image priors for reasoning [22, 3], neglecting important spatiotemporal features.
+This phenomenon is exemplified by a few frames having a dominant influence on response generation.
+To investigate this bias influences hallucination generation in video-based VLLMs, we first utilized a video summarization algorithm [47] to extract the most salient frame $v^{i}$ from $V^{i}$.
+Next, we generated VLLM responses on VidHal using $v^{i}$ instead of $V^{i}$ as the the visual input.
+The influence of image priors is evaluated by identifying overlapping instances where VLLM responses based on the original video $V^{i}$ and the salient frame $v^{i}$ remain consistent, across both correctly and incorrectly predicted orderings.
+As illustrated in Figure 7, our results highlight that VLLMs exhibit strong reliance on image priors in their predictions.
+This phenomenon is particularly severe for smaller models such as VideoLLaMA2 (7B).
+Figure 10: Qualitative examples of VLLM responses on the caption ordering tasks, for the Attribute, Order and Action aspects.
 Figure 8: (Top) Invalid response rates across all models.
@@
 Figure 9: Hallucination misalignment (HM) scores on VidHal, with Random representing HM scores from the random baseline.
-Figure 10: Qualitative examples of VLLM responses on the caption ordering tasks, for the attribute, order and action aspects.
-Image Prior Reliance
-Prior research reveals that VLLMs often leverage image priors for reasoning [23, 3], while ignoring the important spatiotemporal features.
-This phenomenon is exemplified by a few frames having a dominant influence on response generation.
-To investigate whether the image prior issue influences hallucination generation in video-based VLLMs, we first utilized a video summarization algorithm [46] to extract the most salient frame $v^{i}$ from $V^{i}$.
-Next, we generated VLLM responses on VidHal using $v^{i}$ instead of $V^{i}$ as the the visual input.
-The influence of image priors is evaluated by identifying overlapping instances where VLLM responses based on the original video $V^{i}$ and the salient frame $v^{i}$ remain consistent, across both correctly and incorrectly predicted orderings.
-As illustrated in Table 3, our results highlight that VLLMs exhibit strong reliance on image priors in their predictions.
-This phenomenon is particularly severe for smaller models such as VideoLLaMA2 (7B).
-5.4 In-Depth Analysis
-We further leverage three quantitative metrics to identify other deficiencies in VLLM responses.
+5.4 Further Analysis
+Response Stability
+The unique structure of the naive caption ordering task, distinct from existing visual understanding tasks, may lead to inadequate model responses due to potential gaps in understanding.
+To identify these deficiencies, we use the two metrics outlined below:
   • Regurgitation Rate (RR) captures the model’s propensity to consistently generate the same response regardless of input.
 We define this as the maximum proportion of instances in VidHal where a specific caption order is predicted by the VLLM, across all possible orderings.
-  • Hallucination Misalignment (HM) exposes the tendency in visual-language models to favor higher-hallucination captions over lower-hallucination ones in relative caption ranking tasks.
-We introduce two sub-metrics to detect such misalignments: $HM_{3\rightarrow 1}$ and $HM_{3\rightarrow 2}$, representing the proportion of cases where the VLLM selects the most hallucinatory caption over the anchor and the lower-hallucination caption, respectively.
-The general form for both terms can be expressed formally as follows:
-    $HM_{j\rightarrow k}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[\mathcal{Y}^{i,j}_{*}\prec\mathcal{Y}^{i,k}_{*}\right].$ (7)
   • Invalid Response Rate (IRR) measures the proportion of VLLM responses that fail to provide a valid caption order for the naive caption ordering task.
@@
 Despite being able to formulate responses with the correct structure, many VLLMs tend to produce the same caption order regardless of $V^{i}$, as reflected by their high RR scores.
-This is particularly evident in models with limited capacity, such as those equipped with LLMs in the 7B parameter range.
-Advanced VLLMs, such as VideoLLaMA2 (72B), can generally distinguish positive captions from severely hallucinated ones, as reflected by their low $HM_{3\rightarrow 1}$ scores in Figure 9.
-However, all evaluated models possess greater $HM_{3\rightarrow 2}$ scores, indicating difficulty in differentiating between two hallucinatory captions with varying levels of hallucination.
-This gap is especially pronounced in advanced VLLMs like GPT-4o, with $HM_{3\rightarrow 2}$ being significantly closer to the random baseline compared to $HM_{3\rightarrow 1}$.
-These results reinforce that gaps in nuanced video reasoning may potentially drive hallucinatory behavior in VLLMs, a challenge overlooked by existing paired ⟨positive, hallucinatory⟩-based evaluations [29, 50].
+This is particularly evident in models with limited capacity, such as those in the 7B parameter range.
+Hallucination Differentiation Sensitivity
+We investigate the tendency of VLLMs to favor captions with higher hallucination over those with lower degree in the relative caption ranking task.
+For two captions with different hallucination levels $j,k$ where $j>k$, we introduce the following metric to quantify such hallucination misalignment cases:
+$HM_{j\rightarrow k}=\frac{1}{N}\sum^{N}_{i=1}\mathbb{I}\left[\mathcal{Y}^{i,j}_{*}\prec\mathcal{Y}^{i,k}_{*}\right].$ (8)
+which reflects the proportion of cases in which the VLLM selects the caption with a higher level of hallucination $j$ over $k$.
+Specifically, we examine three key cases: when the most hallucinatory caption is chosen over both the lower-hallucination and anchor captions, and when the lower-hallucination caption is selected over the anchor caption.
+These cases are represented by $HM_{3\rightarrow 1}$, $HM_{3\rightarrow 2}$, and $HM_{2\rightarrow 1}$, respectively, with results presented in Figure 9.
+Our findings show that advanced VLLMs, such as VideoLLaMA2 (72B), can generally distinguish positive captions from severely hallucinated ones, as reflected by their low $HM_{3\rightarrow 1}$ scores in Figure 9.
+However, two key observations emerge from our experiments: First, most VLLMs struggle to differentiate the lower hallucinatory caption from the anchor, as evidenced by the gap between $HM_{3\rightarrow 1}$ and $HM_{2\rightarrow 1}$.
+Second, all models exhibit high $HM_{3\rightarrow 2}$ scores, indicating difficulty in distinguishing between two hallucinatory captions with varying degrees.
+These results suggests that gaps in nuanced video reasoning may contribute to hallucinatory behavior in VLLMs, a challenge not addressed by existing ⟨positive, hallucinatory⟩-based evaluation methods.
+[29, 51, 52, 16].
 5.5 Qualitative Results
@@
 Additionally, we propose a novel caption ordering evaluation task to probe the fine-grained video understanding capabilities of VLLMs.
-We conduct extensive experiments on VidHal through the evaluation of eleven VLLMs, exposing their limitations in unexpected hallucination generation.
-Our empirical results shed light on several promising directions for future work: e.g., fine-tuning models with human feedback to reduce undesired regurgitation responses, and countering single-frame priors for improving temporal reasoning.
-These advancements, on the other hand, will mitigate the hallucination problem in video-based VLLMs, thereby improving their robustness for real-world video understanding applications.
-A Benchmark Construction Details
-A.1 Dataset Statistics
-Figure 11 illustrates the distribution of public dataset sources contributing to the visual instances in VidHal.
-Additionally, Figures 12 and 13 depict the distribution of temporal aspects across VidHal and the ground truth answers for the MCQA and caption ordering tasks, respectively.
-One can observe that both temporal aspects and ground truth options are uniformly distributed across our benchmark.
-Figure 11: Distribution of visual instances in VidHal by public dataset source, categorized across the five temporal aspects.
-Figure 12: Distribution of temporal aspects and subaspects in VidHal.
-Figure 13: Distribution of (Left) correct answer options for the MCQA task and (Right) optimal option orders for the caption ordering task.
-A.2 Dataset Development Pipeline
-Figure 14: Specific skills and corresponding questions from the Perception Test dataset chosen for VidHal instance selection, with the matched aspects indicated in brackets.
-Figure 15: Evaluation tasks in MVBench aligned with temporal aspects in VidHal, categorized by aspect.
-Visual Instance Selection
-To ensure a rich coverage of temporal aspects and visual diversity, we methodically selected video instances from four public datasets: TempCompass [38], Perception Test [43], MVBench [28], and AutoEval Video [6].
-Given the unique characteristics of each dataset, we outline the specific guidelines adopted for each dataset below:
-  • TempCompass encompasses five temporal aspects: Action, Speed, Direction, Event Order, and Attribute Change.
-As most of these aspects align with those chosen to construct VidHal, we retain all video instances except those related to speed.
-TempCompass includes four evaluation tasks: MCQA, Yes/No QA, caption matching, and caption generation.
-Given the conciseness of captions in the latter two tasks, their information can often be subsumed within the more detailed QA-based annotations.
-Therefore, we focus exclusively on MCQA and Yes/No QA annotations to create an informative anchor caption.
-  • Perception Test spans various skill and reasoning domains to thoroughly evaluate VLLMs’ perception and understanding abilities.
-Our inspection of these evaluation dimensions reveals alignment between the semantics, physics, and memory skill areas, as well as descriptive and explanatory reasoning dimensions, with the temporal aspects of action, order, and event order.
-Accordingly, we limit our video selection in Perception Test to these specific pillars.
-Additionally, we review the question templates adopted in these areas and select video instances with question-answer pairs that support VidHal’s evaluation objectives.
-The specific skills and associated questions chosen are detailed in Figure 14.
-  • MVBench includes twenty video understanding tasks with question-answer pairs designed to challenge the reasoning capabilities of VLLMs.
-Similar to the Perception Test, we identify the tasks relevant to the temporal aspects in VidHal and focus on collecting videos belonging from these tasks.
-The specific tasks for each aspect are presented in Figure15.
-We observe that MVBench contains repeated use of certain scenarios across tasks, indicated by similar question templates.
-To enhance caption diversity and minimize redundancy, we limit the number of examples for each unique scenario.
-The collected instances cover all five temporal aspects of VidHal.
-  • AutoEval-Video evaluates open-ended response generation in VLLMs through questions with detailed answers across nine skill dimensions.
-We focus on instances related to the state transition area, specifically assessing changes in object and entity attributes.
-For each instance, we retain the only answers to associated questions as they act as informative, long-form captions for the video.
-Incorrect Anchor Captions
-A minority of videos contain anchor captions misaligned with their content, often due to noisy metadata.
-Such discrepancies subsequently lead to undesirable hallucinatory captions.
-To remove such instances, we use BLIP2 [27] to calculate frame-text matching scores across all video frames, selecting the maximum score as the representative video-text alignment score.
-Examples with incorrect anchor captions typically achieve low alignment scores, which are discarded as noisy instances.
-LLM-based Caption Generation
-We utilize GPT-4o’s [41] text processing and generation capabilities to generate an anchor caption for each selected video, based on metadata from its original public dataset source.
-This metadata includes QA-based annotations for TempCompass, Perception Test, and MVBench, along with long-form answers for AutoEval-Video.
-The anchor caption is subsequently used as input for GPT-4o to generate corresponding hallucinatory captions.
-To ensure the generated hallucinatory captions meet high-quality standards, we employ a detailed prompt adopting the following strategies to guide GPT-4o’s output:
-  • Aspect-specific definitions which outline the characteristics of each aspect to be varied, prompting GPT-4o to modify anchor captions accordingly.
-  • Caption construction guidelines that define the structure, format, and hallucination levels required for the generated captions.
-  • In-context examples to illustrate the desired form of each hallucinatory caption for each aspect.
-The prompts for generating anchor and hallucinatory captions are shown in Figures 16 and 17, respectively.
-Definitions for each aspect are provided in Figure 18, and aspect-specific in-context examples are detailed in Figures 19 to 27.
-Separate in-context examples are provided for each Attribute subaspect of Shape, Size, Color, Count, and State Change to account for their distinct natures.
-Figure 16: Prompts used for generating the anchor caption from (Top) QA-based annotations and (Bottom) long-form captions.
-Figure 17: Prompt for generating aspect-specific hallucinatory captions based on anchor captions and in-context examples.
-Figure 18: Definitions incorporated into the prompt for generating hallucinatory captions for each aspect, with separate definitions provided for each sub-aspect in the Attribute aspect.
-Figure 19: In-context examples for Size sub-aspect.
-Figure 20: In-context examples for Shape sub-aspect.
-Figure 21: In-context examples for the Color sub-aspect.
-Figure 22: In-context examples for the Count sub-aspect.
-Figure 23: In-context examples for the State sub-aspect.
-Figure 24: In-context examples for the Object aspect.
-Figure 25: In-context examples for the Event Order aspect.
-Figure 26: In-context examples for the Action aspect.
-Figure 27: In-context examples for the Direction aspect.
-Figure 28: Prompt template for evaluating the quality of generated captions for the GPT-4o, Gemini-1.5 Flash, and LLaMA3 (70B) models.
-Figure 29: Question prompts for evaluating caption quality based on the three assessment criteria.
-Prompts with the placeholder <option> are applied individually to the anchor and hallucinatory captions.
-For question associated with order quality, <option_A> and <option_B> are replaced with the corresponding hallucinatory caption options shown to the LLMs.
-Caption Quality Scoring
-To identify video instances with the high quality generated captions, we utilize powerful LLMs to evaluate the quality of generated captions.
-The captions are assessed is based on three specific criteria:
-  • Realism determines whether generated scenarios are plausible.
-  • Ordering Quality evaluates whether the hallucination level ordering is appropriate.
-  • Relevance ensures that deviations from the anchor caption align with the designated aspect.
-Binary questions are used to evaluate captions for each criterion, assigning a score of 1 for positive responses, i.e., ”yes”, and 0 otherwise.
-The scores for each criterion are averaged across all models and prompts, and then summed across all criteria to produce a final quality assessment score for the generated captions of a video instance.
-We evaluate each set of captions using three LLMs: GPT-4o, Gemini-1.5 Flash [44], and LLaMA3 (70B) [12] along with three variants for each binary question.
-This ensemble of both models and prompts enhances the robustness of our evaluation..
-Figures 28 and 29 provide details of the criterion-specific quality assessment queries and the prompt templates employed for each LLM.
-We select the top 1,000 examples with the highest quality assessment scores to construct VidHal.
-Figure 30: Qualitative examples of video instances and their corresponding generated captions in the VidHal Benchmark, across the five temporal aspects.
-A.3 Additional Dataset Examples
-We provide additional qualitative examples of video instances and their corresponding captions in Figure 30 for each of the five temporal aspects.
-B Human Validation Details
-B.1 Human Validation Process
-As varying hallucination levels are a distinctive feature of our benchmark, we prioritize validating the robustness of caption ordering produced by our annotation pipeline.
-Each anchor caption is derived from the original video metadata, making it the most accurate reflection of the video content.
-Our primary objective is to ensure that the ordering of hallucinatory captions aligns with human judgment.
-To achieve this, human annotators are shown the video instance along with both hallucinatory captions and are tasked with selecting the caption that better aligns with the video content, as illustrated in Figure 31.
-Each video instance is reviewed by multiple annotators, with the final human-aligned order determined through a majority vote and compared with our automatically generated order.
-Figure 31: Pipeline for validating the quality of generated caption orders in VidHal.
-For each instance, human annotators are provided with the video and its associated hallucinatory captions.
-The annotators then select the caption that best aligns with the video content.
-The selected response is subsequently checked for consistency with the caption with lower hallucination according to our annotation process.
-B.2 Misaligned Instances
-Table 4 lists video instances that fail to meet the majority agreement threshold established by our annotation process.
-We additionally provide the corresponding human agreement scores for each instance.
-Video ID Agreement Score
-action_55 0.429
-action_88 0
-action_90 0.308
-action_118 0.200
-action_153 0.250
-order_60 0.500
-order_109 0.154
-attribute_90 0.400
-attribute_180 0.071
-attribute_192 0.188
-object_25 0.375
-object_170 0
-direction_188 0.400
-Table 4: Video examples with generated caption orders misaligned with the human-preferred order during the quality verification process.
-The agreement score indicates the proportion of human respondents who select the ordering generated in our annotations.
-C Evaluation Pipeline Details
-C.1 Evaluation Task Prompts
-Figures 32 and 33 present the prompts used for the MCQA and naive caption ordering tasks, respectively.
-The same prompt used for both the MCQA task and the paired questions in the relative caption ordering task.
-Figure 32: Prompt template for the MCQA and relative caption ordering evaluation tasks.
-Figure 33: Prompt template for the naive caption ordering evaluation task.
-C.2 Relative Order Parsing
-Prompting the VLLM to predict the order of captions based on their hallucinatory level in the relative caption ordering task involves asking a series of paired questions derived from different caption combinations.
-However, providing the model with all possible pairs at once may result in cyclic and non-transitive orderings.
-To address this, we present each caption pair to the VLLM in a systematically selected sequence, beginning with two paired questions.
-The final paired question is presented to the model to resolve inconsistencies if the multiple possible orderings can be derived from the responses to the first two paired questions.
-The responses across all paired questions presented to the VLLM is then parsed according to the workflow illustrated in Figure 34.
-Figure 34: Decision tree for determining the final caption order based on VLLM responses to paired questions in the relative caption ordering evaluation task.
+We conduct extensive experiments on VidHal through the evaluation of thirteen VLLMs, exposing their limitations in unexpected hallucination generation.
+Our empirical results shed light on several promising directions for future work: e.g., incorporating a broader range of temporal features during pretraining and mitigating single-frame priors to enhance temporal reasoning.
+These advancements will help to address the hallucination problem in video-based VLLMs, enhancing their robustness for real-world video understanding applications.